OpenStudy (anonymous):

-2, -4 , -8 , -16... find a 11

OpenStudy (diyadiya):

\[\Large a_n=a_1+(n-1)d\] \[\Large a_{11}=a_1+(11-1)d\] To find d: \( d=a_2 - a_1 \) Here \(a_1= -2 \) \(a_2 = -4 \) \(d=-4 - (-2)=-4+2= -2 \) Now plug in the values of \(a_1 \ and \ d \) in the equation & find \(\Large a_{11}\)